• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:02:26 on June14,2004

    Entry number 125736

    keyword=day shift summary

    Day shift summary:
    
    8:00 shift started with no beam. Accelerator recovering from a main power breaker failure.
    9:00 oppotunistic controlled access for detector, target and spectrometer work.
    10:50 MCC caled, they are ready to send tuning beam. 
    11:07 -Run Coordinator recommandations:
    when beam comes back one should ask that MCC saves the beamline and loads the new file that J. Benesh has produced (elog - entry 1218415) and run spot++ to check that the raster is effectively 4*4
    11:10: Left Q1 tripped and reset.
    11:20 Back to beam permit. Moved target to BeO for beam tuning.
    -12:25 MCC tuned beam with old optics and saved. Then loaded new optics, Tuned beam. We did raster check.
    12:30 Moved to 4He target, started target tests.
     1x1 mm raster, 28 uA.
    Fan speed 24 Hz., 40, 50, 60 and 70 Hz.
    Then repeat with raster=3.5x3.6mm (max).
    13:40-14:20 test with target operating temperature vs cooling power.
    6.6K to 7.6K give extra 115W HP heater.
    14:35-14:51 dithering run 2125
    Get back to production run
    15:00 run 2126, Compton is tuning while the run is going on. position
    feedback was off. Then it turned back on.
    15:55 Compton tuning finished. End run 2126.
    15:55 Starting run 2127, still going on at the end of the shift.